WACC Midwest Electric Company (MEC) uses only debt and common equity. It can borrow unlimited amounts at an interest rate of rd-1196 as long as it finances at its target capital structure, which calls for 30% debt and 70% common equity. Its last dividend (D0) was $2.30, its expected constant growth rate is 4%, and its common stock sells for $21. MECs tax rate is 40%. Two projects are available. Project A has a rate of return of 14%, while Project Bs return is 10%. These two projects are equally risky and about as risky as the firms existing assets a. What is its cost of common equity? Round your answer to two decimal places. b. What is the WACC? Round your answer to two decimal places.
Answer #1

As per dividend discount model,

P0 = D0(1+g) / ke - g, we have to calculate ke or cost of common equity

21=2.3(1.04) / ke - 0.04

ke = 15.39%

a. Cost of common equity is 15.39%

b. WACC = ke * Wa + kd * Wb

=15.39% * 70% + 11(1-0.40) * 30%

=12.75%

WACC is 12.75%